OpenStudy (anonymous):

Find the domain of inverse sine of (x^2-1)

OpenStudy (anonymous):



OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

close. x can be negative

OpenStudy (anonymous):

Since the domain of \(\sin^{-1}(y)\) would be \(-1\leq y\leq 1\) then we would say \[ -1\leq x^2 - 1\leq 1 \]And solve for \(x\).

OpenStudy (anonymous):

so what would be the result?

OpenStudy (anonymous):

\[ 0\leq x^2\leq 2 \]\[ 0\leq |x| \leq \sqrt 2 \]Finally \[ -\sqrt 2\leq x \leq \sqrt 2 \]

OpenStudy (anonymous):

why is it -root 2

OpenStudy (anonymous):

\(\sqrt{x^2} = |x|\) and \(|x|\leq a \implies -a\leq x \leq a\)

OpenStudy (anonymous):

ok, thanks!